OpenStudy (anonymous):

The length of a rectangle is 5 more than 3 times its width. The perimeter of the rectangle is 58 inches. What is the length of the rectangle?

OpenStudy (anonymous):

A) 23 in. B) 29 in. C) 32 in. D) 35 in.

OpenStudy (anonymous):

L = Length W = Width P = Perimetre P = 2(L + W) L = 3W + 5 L + W = 41 3W + 5 + W = 41 4W = 36 [ W = 9 inches ] L = 3W + 5 L = 3(9) + 5 [ L = 32 inches ] Hope this helps

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

your welcome I really do hope u get it right

OpenStudy (mathmate):

@LesTwins It would be nice to do a check when solving system of equations For example, with the solution W=9, L=32, Perimeter= 2(W+L) = 2(9+32) = 82 \(\ne\) 58 as given in the question. So a review of the solution is necessary.

OpenStudy (mathmate):

Also, most of the time, we can do a substitution in situ to avoid a system of equations of two unknowns, for example: Given: Perimeter = 58 inches and Length = 3*width +5 inches. Let W=width, L=length, then L=3W+5 Perimeter = 2(W+L)= 2(W+3W+5) = 58 Divide by two on both sides 4W+5=29 4W=24 W=6 and L=3W+5 = 23 Check: Perimeter = 2(W+L)=2(6+23)=2(29)=58 inches as given. Answer: Width = 6 inches, Length = 23 inches.